40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Aug 18 18:33:57 2014, ericq, Update, Green Locking, Yarm Green PDH yLoop.pdfcrap.pdf
    Reply  Tue Aug 19 01:01:36 2014, Jenne, Rana, Update, Green Locking, Yarm Green PDH YPDH_OLG.pdfYPDH_noise.pdf
       Reply  Tue Aug 19 18:32:40 2014, ericq, Update, Green Locking, Yarm Green PDH 
       Reply  Tue Aug 19 23:11:15 2014, Jenne, Update, Green Locking, Yarm Green PDH ErrAndCtrlSpectra_VoltsPerRtHz.png
       Reply  Wed Aug 20 02:38:41 2014, Jenne, Update, Green Locking, Yarm Green PDH - requirement NoiseConsideration.pdf
          Reply  Wed Aug 20 18:05:18 2014, Jenne, Update, Green Locking, Yarm Green PDH - requirement YPDH_noise.pdf
       Reply  Wed Aug 20 04:09:21 2014, ericq, Update, Green Locking, Xarm Green PDH Xbodes.pdfXspectra.pdfscopeSweep.jpg
          Reply  Wed Aug 20 16:10:43 2014, ericq, Update, Green Locking, Xarm Green PDH XspectraCombined.pdf
             Reply  Wed Aug 20 21:09:16 2014, ericq, Update, Green Locking, Xarm Green PDH scopeSweep.jpgXspectra.pdfYspectra.pdfXYcomp.pdf
Message ID: 10416     Entry time: Wed Aug 20 18:05:18 2014     In reply to: 10412
Author: Jenne 
Type: Update 
Category: Green Locking 
Subject: Yarm Green PDH - requirement 

 

 I calibrated the control signal from Volts to Hz using the rough PZT calibration of 5MHz/V for the Yend NPRO.  

For the error signal, Q said that the Yarm PDH peak-to-peak height was about a factor of 100 smaller than the Xarm, so I used a calibration of 1.9e7 Hz / V.

Then, from Q's Mist simulation including the high Xarm loss, and the plot that he posted in the control room, the CARM linewidth looks like it is about 2pm.  This is the number that I have included on today's plot.  Note though that yesterday I was using a linewidth of about 30pm, which I got from an Optical simulation about a year ago.  I do not know why these numbers come out an order of magnitude different!      The CARM linewidth is actually about 20 pm.  Both Q and I failed at reading log-x plots yesterday.  I have corrected this, and replotted.

Anyhow, here's the Yarm noise spectra calibrated plot:

YPDH_noise.pdf

I have emailed Kiwamu, but haven't heard back from him yet on what the original design considerations were, if he remembered us ever using a boost, etc.  What this looks like to me is that we need to do some serious work to get the noise down.  Maybe fixing the gain peaking and triggering the boost will get us most of the way there?

ELOG V3.1.3-